OpenStudy (anonymous):

PLEASE HELP!!! If f(x) = x - 4, what does f(3) represent? The value of (x - 4) when x = 3 The value if x when (x - 4) = 3 The value of 3f(x) when x = 3 The value of f(x) when (x - 4) = 3

OpenStudy (solomonzelman):

plug in 3 for x. f(3) = (3)-4 f(3) = ?

OpenStudy (anonymous):

f(3) = -1?

OpenStudy (solomonzelman):

Yes , there you go !

OpenStudy (anonymous):

But that isn't an option :( I'm confused.

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

Then check your question

OpenStudy (anonymous):

ok. thanks :)

OpenStudy (solomonzelman):

I see...

OpenStudy (solomonzelman):

Well which of the choices below A) The value of (x - 4) when x = 3 B) The value if x when (x - 4) = 3 C) The value of 3f(x) when x = 3 D) The value of f(x) when (x - 4) = 3 represents "plug in 3 for x into "x-4" ?

OpenStudy (anonymous):

A?

OpenStudy (solomonzelman):

Yeah, !

OpenStudy (anonymous):

Thanks :)

OpenStudy (solomonzelman):

You are a pleasure to help to ! You welcome !